### 25 Jan 2024 v1

# Revealing hidden structures and symmetries in nonequilibrium transport

Abstract
Recent results have shown how to partition the parameter space of Markov systems into dynamical equivalence classes. These equivalence classes structure transport properties in such a way that makes, among other features, their responses fully symmetric. In this note, I illustrate this approach on two representative systems. First, I derive analytical expressions for the equivalence classes of a disordered ring model. Second, I verify on a model of ion transport that, within an equivalence class, the response of coupled currents is symmetric both near and far from equilibrium.
Context and objectives
Nonequilibrium transport is traditionally characterized by quantities such as currents, affinities or the entropy production of the system. Current fluctuations and related fluctuation theorems also provide additional constraints and insights on the behavior of nonequilibrium systems.
However, despite their relevance, these core nonequilibrium concepts are not enough to build a well-defined nonequilibrium theory. What do I mean by that?
Take a nonequilibrium system of your choice. It is, by definition, out of equilibrium - but with respect to what? There is a continuum of equilibrium states that could serve as a reference point. And without a clearly defined equilibrium state, we cannot meaningfully characterize the nonequilibrium response of the system as there is no unique way to vary the set of parameters.
Choosing an arbitrary reference equilibrium state doesn’t solve the issue either. From a given equilibrium state, there is an infinite number of ways you can drive the system out of equilibrium. While in the linear response regime it doesn’t really matter since all response coefficients are identical regardless of the nonequilibrium paths you choose, this is no longer true in the nonlinear regime [1].
So, how do we remove this arbitrariness and connect equilibrium and nonequilibrium dynamics in a way that is:
  • Well-defined, i.e. there is a unique way to vary a system’s parameters to generate all possible nonequilibrium conditions
  • Intrinsic, i.e. the association emerges from the dynamics of the system itself.
To structure the space of possible dynamics while satisfying these requirements, I previously introduced the concept of dynamical equivalence classes [2, 3]. In addition to offering a well-defined nonequilibrium theory, they also bring extra benefits. For example, the current fluctuations can be obtained from an equilibrium dynamics and their nonlinear responses takes an especially simple form [4]. These results go beyond the traditional current fluctuation theorem [4, 2].
This new framework thus offers a new lens to understand nonequilibrium systems. Yet, it remains difficult to form a physical intuition around the equivalence classes themselves. This motivated my recent analysis of a molecular motor toy model, for which I derived exact expressions for its equivalence classes and transport properties [5].
In this note, my objective is to provide additional examples of how this framework plays out for specific systems. I first provide an analytical expression for the equivalence classes of a disordered ring model. Next, using a model of ion transport in membranes, I illustrate the symmetries of the response coefficients when multiple currents are present. I hope these examples will help to build physical intuition and kickstart the broader application of this framework, e.g. to the design and control of nonequilibrium systems.
I Transport on a Disordered Ring
Let’s consider a $N$-state system characterized by the allowed transitions $i\rightarrow i\pm 1$ with periodic boundary conditions:
$P$ $=$ $\begin{pmatrix}0&p_{1}&&&&1-p_{1}\\ 1-p_{2}&0&p_{2}&&&\\ &1-p_{3}&0&p_{3}&&\\ &&\ddots&0&\ddots&\\ &&&1-p_{N-1}&0&p_{N-1}\\ p_{N}&&&&1-p_{N}&0\end{pmatrix}\,.$ (7)
The space of possible dynamics constitues a $N$-dimensional cube $\Sigma=\{0<p_{i}\leq 1\}$ with $i=1,\cdots,N$. Within this space, the set of equilibrium dynamics forms an $(N-1)$-dimensional hypersurface defined by
$\prod_{i=1}^{N}\frac{p_{i}}{(1-p_{i})}=1\,.$ (8)
Outside of this hypersurface, the affinity
$A=\ln\prod_{i=1}^{N}\frac{p_{i}}{(1-p_{i})}$ (9)
differs from zero and the system is out of equilibrium. From this expression, it is already clear that there is no unique way to define a nonequilibrium response. There is an infinite numbers of paths in $\Sigma$ with the same affinities (and starting from the same equilibrium state) but with different currents [5].
What we can do instead is to partition the space $\Sigma$ into dynamical equivalent classes [2]. Two dynamics $P$ and $Q$ belong to the same class if there exists a scalar $\rho$ such that:
$P\sim Q\quad{\rm if}\quad P\circ P^{T}=\rho\,\,Q\circ Q^{T}$ (10)
where $\circ$ denotes the Hadamard product.
We can use relation (10) to generate the equivalence classes of the disordered ring (7). Let’s consider an equilibrium dynamics $Q=\bar{Q}$ with transition probabilities $\bar{q}_{i}$ satisfying (8). Any dynamics $P$ in $[Q]$ then satisfies the system of $N$ polynomial equations
$p_{i}(1-p_{i+1})=\rho\,\bar{q}_{i}(1-\bar{q}_{i+1})$ (11)
for some value of the parameter $\rho$.
This polynomial system can be solved using Gröbner bases. For $N=3$, the solution of (11) takes a simple form. Introducing the quantities
$a_{i}(\rho)\equiv\rho\,\bar{q}_{i}(1-\bar{q}_{i+1})-1\,,$ (12)
the equivalence class $[\bar{Q}]$ is given by the set of dynamics
$p^{\pm}_{i}(\rho)=\frac{1}{2a_{i+1}}(\pm\sqrt{\Delta}+a_{i-1}-a_{i}+a_{i+1})$ (13)
with
$\Delta=(a_{1}^{2}+a_{2}^{2}+a_{3}^{2})-2(a_{1}a_{2}+a_{2}a_{3}+a_{1}a_{3})-4\,a_{1}a_{2}a_{3}\,,$ (14)
and where $\rho$ varies between $[0,1]$ and the coefficients $a_{i}$ satisfy periodic boundary conditions, i.e. $a_{N+1}=a_{1}$ and $a_{0}=a_{N}$. In this parametrization, equilibrium occurs when $\rho=1$ or equivalently $\Delta=0$ [6]. For $\rho<1$ ($\Delta\neq 0$), both solutions $p^{\pm}_{i}$ belong to $[\bar{P}]$. They are time-reversed of each other, with opposite affinities and currents.
Figure 1: The space $\Sigma$ for $N=3$ and its decomposition into equivalence classes. Each curve corresponds to a different equivalence class $[\bar{Q}$]. Each class crosses the equilibrium hypersurface once and spans the full range of affinities.
We thus see that dynamical equivalence classes have a clear physical interpretation. For the disordered ring (7), each equivalence class determines a curve in $\Sigma$ that crosses the equilibrium hypersurface only once and whose affinities span the entire range [,][$\infty$,$\infty$][ - ∞ , ∞ ] as the curves tends to $p_{i}=0$ or $p_{i}=1$ for all $i$ (Figure 1).
As a result, within an equivalence class the nonequilibrium behavior is well-defined: there is a unique equilibrium state [10], all possible affinities are spanned, and there is a unique way to vary the parameters while driving the system out of equilibrium. Moreover, for each dynamics in $[\bar{Q}]$ its time-reversed one also belongs to $[\bar{Q}]$ with opposite affinities and currents. These constitute general properties of dynamical equivalence classes [4, 5].
Now that we understand the structure of equivalence classes and their physical interpretation, we can use them to study the nonequilibrium behavior of the disordered ring system.
[…]
The higher-order cumulants $K^{(n)}(A)$ can then be directly obtained from the mean current as
$K^{(n)}(A)=\frac{{\rm d}J^{n}(A)}{{\rm d}A}\,.$ (15)
Note that, within an equivalence class, this relation is valid outside the linear regime and at all orders $n$. In contrast, the traditional fluctuation relations take a more complex forms [4].
[…]
The solution (14) of the disordered ring offers a window into the nonequilibrium dynamics of systems with one independent current. For systems for more than one currents, additional symmetries emerge within an equivalence class [4]. We illustrate these additional symmetries in the next section.
II Transport of ions through a membrane
Let’s now turn to a model of ions transport as an archetype for systems with two independent currents. Following Hill [8], consider a cell surrounded by a membrane that separates the cell’s interior (In) from its environment (Out). A complex E, which can exist in two conformations E and E*, has binding sites for ions L and M. These sites are accessible to inside molecules in configuration E only, and to outside molecules in configuration E* only. L can be bound only when M is already bound on its site (Figure 3a).
Figure 3: Ion transport model with two independent currents. (a) Mechanism for the transport of ions M and L across the membrane. In the normal mode of operation, molecule M has a larger concentration inside than outside, $[{\rm M_{i}}]>[{\rm M_{o}}]$, while the opposite holds for molecule L, $[{\rm L_{o}}]>[{\rm L_{i}}]$. The complex E acts as a free energy transducer and utilizes the M concentration gradient to drive molecules of L from inside to outside against its concentration gradient. For example, in the case of the Na/K-ATPase complex M and L would correspond to ${\rm K}^{+}$ and ${\rm Na}^{+}$, and transport would be coupled to ATP consumption. (b) Kinetic diagram. (c) Cycle decomposition. The cycles $a$ and $b$ are chosen as the two independent cycles. The positive orientation is chosen counterclockwise (adapted from Hill [8]).
The system has $N=6$ states connected by the kinetic diagram depicted in Figure 3b. The parameter space $\Sigma$ is $8$-dimensional while the set of equilibrium dynamics form an $6$-dimensional manifold in $\Sigma$. Each equivalence class corresponds to a $2$-dimensional manifold [13, 14].
Transport is determined by two independent currents, here measured along cycles $a$ and $b$ (Figure 3c). In Figure (4), we depict the currents $J_{a}(\boldsymbol{A})$ and $J_{b}(\boldsymbol{A})$ as a function of the affinities $A_{a}$ and $A_{b}$ in a given equivalence class. As expected, the currents are coupled, i.e. a change in $A_{a}$ will drive the current $J_{b}$ and vice versa. They display a nonlinear behavior and, beyond the anti-symmetry $\boldsymbol{J}(-\boldsymbol{A})=-\boldsymbol{J}(\boldsymbol{A})$ characteristic of equivalence classes, no other symmetry is apparent.
Figure 4: Coupled currents of the ion transport model. Currents $J_{a}$ and $J_{b}$ and their isolines (solid curves) as a function of the affinities $A_{a}$ and $A_{b}$ in a given equivalence class. The equivalence class is defined by the equilibrium dynamics $p_{12}=0.54,p_{16}=0.46,p_{21}=0.9,p_{23}=0.04,p_{25}=0.06,p_{32}=0.06,p_{34}=0.94,p_{43}=0.65,p_{45}=0.35,p_{52}=0.06,p_{54}=0.38,p_{56}=0.56,p_{61}=0.58,$ and $p_{16}=0.42$. The affinities $\boldsymbol{A}$ take values in the range $[-10,+10]$.
Remarkably, despite the nonlinear behavior of the currents, their response is fully symmetric [5] both near and far from equilibrium:
$\frac{\partial J_{a}}{\partial A_{b}}(\boldsymbol{A})=\frac{\partial J_{b}}{\partial A_{a}}(\boldsymbol{A})$ (16)
for all values of the affinities $\boldsymbol{A}$ in a given equivalence class (Figure 5). The symmetry (16) is also confirmed for all other equivalence classes (a Scilab code to generate the response coefficients for an arbitrary equivalence class is available upon request). Higher-order fluctuations can be obtained by further derivation of the currents and their responses are also symmetric [5].
Figure 5: The currents response is symmetric both in the linear and nonlinear regimes. The response coefficients $\partial\boldsymbol{J}/\partial\boldsymbol{A}$ display a complex structure in the nonlinear regime. Even so, the cross-response coefficients are identical both near and far from equilibrium. Parameters take the same values as in Fig. (4).
III Outlook
Properties like (15) and (16) confirm the importance of dynamical equivalence classes as a central concept in nonequilibrium thermodynamics. Where do these remarkable properties come from?
In the traditional way to look at nonequilibrium systems, affinities are varied along an arbitrary path in the parameter space (typically by keeping some parameters constant while varying the others). This makes every dynamics along the path unrelated to each other and with possibly very different dynamical properties. In contrast, within an equivalence class, all systems share common dynamical properties through relation (10). As a result, within an equivalence class the current generating functions of the different dynamics all have the exact same shapes, modulo a translation [2, 3]. All the fluctuations and reponse coefficients are thus directly related to each other, and equilibrium and nonequilibrium fluctuations are essentially identical.
Dynamical equivalence classes thus offer new insights by providing a well-defined structure and additional symmetries for nonequilibrium systems. Beyond their theoretical interests, they bring intriguing perspectives for the design and control of nonequilibrium systems.
Disclaimer. This paper is not intended for journal publication.